Suppose that we have a sample space with seven equally likely experimental outcomes: E1, E2, Es, EA, E5, E, and E. Let a. Finb. Find P(AUB) (to 2 decimals) Are A and B mutually exclusive? -Select your answer- c. Find Ac, Ce, P(A), and P Ce) A -Selectd. Find AU Be and P(AUBe). Select your answer- (to 2 decimals). e. Find P(BUC) (to 2 decimals)
Answer #1

a. P(A) = 2/7 = .29, P(B) = 2/7 = .29,P(C)= 3/7 =.43

b.P(A U B) = P(A) + P(B) = .29+.29= .58

A and B are mutually exclusive

c. Ac = [E3, E4, E5, E6, E7] answer - iii

Cc= [E1, E4,E6,E7] answer - V

P(Ac) = 1 - P(A) = 1-.29 = .71

P(Cc) = 1 - P(C) = 1- .43 = .57

d. A U Bc

here A =[E1, E2], Bc = [E1,E2, E5, E6, E7]

A U Bc = [E1,E2, E5, E6, E7] answer - i

P (A U Bc) = P(A) - P(A n B)

here A and B are mutually exclusive So A n B = 0

P(A U Bc) = .29-0 = .29